A minor league baseball team plays 90 games in a season. if the team won 14 more that three times as many gmaes as they lost, how many wins and losses did the team have?
Answer by macston(5194)

.
L=number lost; W=games won=3L+14
.
Without ties:
L+W=90
(L)+(3L+14)=90
4L+14=90
4L=76
L=19
.
W=3L+14=3(19)+14=57+14=71
.
ANSWER: They lost 19 and won 71.
.
CHECK:
W+L=90
71+19=90
90=90
.